A girl is sledding down a slope that is inclined at 30º with respect to the horizontal. The wind is aiding the motion by providing a steady force of 195 N that is parallel to the motion of the sled. The combined mass of the girl and the sled is 73.2 kg, and the coefficient of kinetic friction between the snow and the runners of the sled is 0.226. How much time is required for the sled to travel down a 212-m slope, starting from rest?
Answer #1

Acceleration a = Fnet/m = [mg sin theta +F - us mg cos theta]/m

= [73.2*9.8* sin 30 degree +195-0.226*73.2*9.8*cos 30 degree]/73.2

= 5.646 m/s^2

Using second equation of motion along incline, s = 0.5at^2

t = sqrt(2s/a) = sqrt(2*212/5.646)

= 8.666 s answer